Select the set of equations that represents the following situation: The sum of two numbers is 36, and their difference is three more than the smaller number. What are the two numbers?
A) x + y = 36; y – x = x + 3
B) 2y = 36; 36 = x – 3
C) x + y = 36; x – y = 3 – x
D) 2x = 36; 36 = x + 3

Answer by sudhanshu_kmr(1152)


ANS: A) x + y = 36; y – x = x + 3